# RIS-Empowered OTFS Modulation With Faster-than-Nyquist Signaling in High-Mobility Wireless Communications

### v1

@@ V Simulation Results
Fig. 9 evaluates the spectral efficiency of the RIS-OTFS-FTN scheme under various configurations, including same $Q=32$, different $\alpha$, $M$, and $N$, as given in tittle and legend of figure. In (a), with fixed $M=16$ and $N=16$, the spectral efficiency is compared for $\alpha=1,0.9,0.8,0.7$ across $Q=30,20,10$. As $\alpha$ decreases, the spectral efficiency exhibits a significant increase, consistent with the inverse proportionality to $\alpha$ in the normalization factor of (39). This enhancement stems from FTN’s ability to pack more symbols within the same frame duration, as illustrated in Fig. 3. Notably, as shown in Fig. 8, while lower $\alpha$ values degrade BER due to heightened ISI, the spectral efficiency gains are substantial and continue to grow with diminishing $\alpha$, highlighting a clear trade-off between error performance and throughput. Furthermore, increasing $Q$ amplifies the spectral efficiency across all $\alpha$, attributable to the RIS-induced boost in effective channel gain, particularly at higher SNR where the scheme approaches capacity. (b) fixes $Q=10$ and varies $\alpha=1,0.9,0.8,0.7$ with $M=N=8,16,32$. Similar to (a), spectral efficiency increases with decreasing $\alpha$, reinforcing FTN’s efficiency benefits. Intriguingly, smaller grid sizes in $MN$ yield higher spectral efficiency, as the normalization in (39) scales inversely with $MN$, implying that compact frames benefit more from FTN compression relative to the occupied time-frequency resources. In summary, the RIS-OTFS-FTN scheme embodies a strategic trade-off in high-mobility wireless environments: FTN enhances SE at the cost of BER degradation, RIS scaling bolsters both reliability and efficiency, and grid sizing influences the normalized performance. By considering these effects, the proposed framework offers a versatile alternative to conventional OTFS, adaptable to diverse scenarios such as spectrum-constrained urban vehicular networks or energy-limited high-speed links, thereby providing an additional tool to combat challenging propagation conditions.
@@
Figure 8: BER of the RIS-OTFS-FTN scheme with/without RIS or FTN, as well different $M$ and $N$.
@@
Figure 9: Spectral efficiency of the RIS-OTFS-FTN scheme with different cases in (a) $\alpha=1,0.9,0.8,0.7$, $Q=30,20,10$, $M=16$, and $N=16$; (b) $\alpha=1,0.9,0.8,0.7$, $L=10$, $M=8,16,32$, and $N=8,16,32$.

### v2

@@ V Simulation Results
Fig. 9 evaluates the spectral efficiency of the RIS-OTFS-FTN scheme under various configurations, including same $Q=32$, different $\alpha$, $M$, and $N$, as given in title and legend of figure. In (a), with fixed $M=16$ and $N=16$, the spectral efficiency is compared for $\alpha=1,0.9,0.8,0.7$ across $Q=30,20,10,0$. As $\alpha$ decreases, the spectral efficiency exhibits a significant increase, consistent with the inverse proportionality to $\alpha$ in the normalization factor of (40). This enhancement stems from FTN’s ability to pack more symbols within the same frame duration, as illustrated in Fig. 3. Notably, as shown in Fig. 8, while lower $\alpha$ values degrade BER due to heightened ISI, the spectral efficiency gains are substantial and continue to grow with diminishing $\alpha$, highlighting a clear trade-off between error performance and throughput. Furthermore, increasing $Q$ amplifies the spectral efficiency across all $\alpha$, attributable to the RIS-induced boost in effective channel gain, particularly at higher SNR where the scheme approaches capacity. (b) fixes $Q=10$ and varies $\alpha=1,0.9,0.8,0.7$ with $M=N=8,16,32$. Similar to (a), spectral efficiency increases with decreasing $\alpha$, reinforcing FTN’s efficiency benefits. Intriguingly, smaller grid sizes in $MN$ yield higher spectral efficiency, as the normalization in (40) scales inversely with $MN$, implying that compact frames benefit more from FTN compression relative to the occupied time-frequency resources. In summary, the RIS-OTFS-FTN scheme embodies a strategic trade-off in high-mobility wireless environments: FTN enhances SE at the cost of BER degradation, RIS scaling bolsters both reliability and efficiency, and grid sizing influences the normalized performance. By considering these effects, the proposed framework offers a versatile alternative to conventional OTFS, adaptable to diverse scenarios such as spectrum-constrained urban vehicular networks or energy-limited high-speed links, thereby providing an additional tool to combat challenging propagation conditions.
Fig. 10 depicts the BER performance under imperfect CSI with varying estimation error variances $\sigma_{e}^{2}\in\{0,0.1,0.2\}$, and investigates the impact of finite phase resolution on the BER performance. It is observed that while the performance naturally degrades as the estimation accuracy decreases, the system does not suffer from catastrophic failure. Even with a relatively large error variance of $\sigma_{e}^{2}=0.2$, the proposed RIS phase alignment strategy maintains a valid gain, exhibiting a graceful performance degradation. This confirms the robustness of the proposed scheme against system uncertainties in high-mobility scenarios. (b) evaluates the impact of phase quantization resolution $B\in\{1,2,3\}$ bits compared to the ideal continuous phase case. The results demonstrate that the performance gap between the $B=3$ bits quantization and the continuous phase is negligible. Also, the $B=2$ bits configuration incurs only a marginal SNR penalty, approximately 1-2 dB at $\text{BER}=10^{-2}$, while significantly reducing hardware complexity. However, the 1-bit quantization results in a noticeable performance loss due to the large phase mismatch. These results indicate that a low-resolution RIS architecture is sufficient to realize the benefits of the proposed RIS-OTFS-FTN scheme, making it a cost-effective solution for practical deployment.
Fig. 11 illustrates the BER performance as a function of user velocity ranging from 0 to $6000$ km/h, with a fixed high SNR of $20$ dB to isolate the impact of Doppler-induced impairments. The ”operating limit” is defined as the maximum velocity. For the configuration with $M=N=16$ and $Q=16$, the system exhibits robust performance up to approximately $1,700$ km/h. This indicates that for standard high-speed vehicular or railway scenarios (<500<500 km/h), the proposed scheme operates well within its safety margin with negligible Doppler-induced degradation. Increasing the OTFS grid size to $M=N=32$ significantly extends the tolerance limit to roughly $2,800$ km/h. The finer resolution in the Delay-Doppler domain allows for more accurate channel resolvability and effective separation of high-Doppler paths. Further increasing the number of RIS elements to $Q=32$ pushes the operating limit beyond $4,800$ km/h. The enhanced passive beamforming gain provided by the larger RIS array effectively compensates for the SNR degradation caused by severe inter-carrier interference at hypersonic speeds. Increasing the OTFS grid size to $M=N=32$ significantly extends the tolerance limit to roughly $2,800$ km/h. Besides, A larger $N$ reduces the Doppler bin spacing ($\Delta\nu=\frac{1}{NT}$), enabling the receiver to resolve closely spaced high-Doppler paths that would otherwise merge or cause severe inter-carrier interference (ICI) in a coarser grid, which is why the BER gets better with increase of $MN$. In summary, the proposed architecture demonstrates exceptional robustness, making it a viable candidate not only for terrestrial high-mobility networks but also for future non-terrestrial networks (NTN) such as LEO satellite communications.
@@
Figure 8: BER of the RIS-OTFS-FTN scheme with/without RIS, FTN, MP detector, as well different $M$ and $N$.
@@
Figure 9: Spectral efficiency of the RIS-OTFS-FTN scheme with different cases in (a) $\alpha=1,0.9,0.8,0.7$, $Q=30,20,10,0$, $M=16$, and $N=16$; (b) $\alpha=1,0.9,0.8,0.7$, $Q=10$, $M=8,16,32$, and $N=8,16,32$.
Figure 10: BER of the RIS-OTFS-FTN scheme with imperfect CSI in different $\delta_{e^{2}}^{2}$ and varing number of quantized adjusted phases $B$.
Figure 11: Operating limits analysis of the RIS-FTN-OTFS: BER versus User Velocity (km/h) at SNR = 20 dB.